OpenStudy (anonymous):

An object is dropped from a platform 100 feet high. Ignoring wind resistance, what will its speed be when it reaches the ground? _____ ft/s

OpenStudy (zephyr141):

of these four which has the variables for the values you have?\[y-y_{0}=v_{0}t+\frac{1}{2}at^2\] \[v=v_{0}+at\] \[v^2=v_{0}^2+2a \Delta y\] \[\Delta y=\frac{v_{0}+v}{2}t\] also remember that \[\Delta y= y-y_{0}=y_{final}-y_{initial}\]

OpenStudy (zephyr141):

once you know which to use it's as easy as plug and chug